What is x−y=−5 in slope intercept form
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

start with x - y = -5
subtract x from both sides of the equation to get -y = -x - 5
multiply both sides of the equation by -1 to get y = x + 5
the general form of the slope intercept form of the equation for a straight line is y = mx + b.
m is the slope.
b is the y-intercept.
in this problem, m = 1 is the slope and b = 5 is the y-intercept.
the slope intercept form of this equation is y = x + 5.
the graph of the equation is shown below.
both forms of the equation are shown.
they draw the same line since they are equivalent.

two points on the line are used to find the slope.
the points are (0,5) and (5,10).
slope is (y2 - y1) / (x2 - x1)
assign (5,10) to (x2,y2) and assign (0,5) to (x1,y1).
slope = (y2-y1)/(x2-x1) = (10-5)/(5-0) = 5/5 = 1.

the y-intercept is the value of y when x is 0.
to find, set x to 0 and solve for y.
y = x + 5 becomes y = 5.
the y-intercept is 5 at the coordinate point of (0,5).

here's the graph.


Answer by ikleyn(53751)

.
What is x-y = −5 in slope intercept form
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The slope-intercept form of linear equation is y = mx + b.


To get this form, variable "y" should be expressed via variable "x".


In your case, to get this form, simply move "y" to the right 
and move 5 to the left. Do not forget to change the signs of these terms on the way.


You will get then

    x + 5 = y.


Now swap left side and right side

    y = x + 5.


It is what you want to get.


ANSWER.  The slope-intercept form of the given equation is y = x + 5.

Solved.

---------------

In the slope-intercept form,  " m "  is the slope.

In your case,  the slope is  m = 1.  You can see it directly and explicitly from the slope-intercept equation - - -
- - - therefore,  there is  NO  NEED  to re-calculate the slope again,  as the other tutor does.